10x-0.62=9(x-0.95)
Answer by checkley71(8403)

10X-.62=9(X-.95)
10X-.62=9X-8.55
10X-9X=-8.55+.62
X=-7.93
PROOF
10*-7.93-.62=9(-7.93-.95)
-79.3-.62=9(8.88)
-79.92=79.92